A child and sled with a combined mass of 2.0 kg slA child and sled with a combined mass of 52.0 kg slide down a frictionless slope. If the sled starts from rest and has a speed of 3.70 m/s at the bottom, what is the height of the hill?
Answer #1

Ans=   mgh = (1/2)(m)(v^2)

g*h = 1/2* v^2

9.81*h=0.5*3.70^2

h=0.697 m